What is the resistance of a toaster that draws a current of 12A when connected to a 120V circuit?
1
Expert's answer
2019-11-21T10:02:06-0500


We depict the equivalent equivalent circuit

We write Ohm's law

"I=\frac{U}{R}"

Where do we write

"R=\frac{U}{I}=\frac{120(V)}{12(A)}=10(Om)"